# Unveiling the Potential of Multimodal Retrieval Augmented Generation with Planning

### v1

@@ 4.2.2. Retrieval and Generation
At each iterative stage, the selection of retrieval actions determines whether or not and which retrieval API is invoked, either text retrieval or image retrieval. We leverage Google Web Search and Google Image Search as our primary retrieval API service. For each text retrieval request, we retrieve the top-k search results to ensure that only the most relevant information is retained. These results undergo preprocessing through the Jina API framework  [^2] https://github.com/jina-ai/reader, which transforms the raw web content into structured representations better suited for MLLM consumption. The visual retrieval pipeline captures full-page screenshots of search results, employing a set of systematic human-crafted rules processes to eliminate extraneous elements such as white space and original query images. To maintain computational efficiency while preserving result quality, we restrict image retrieval to 3-6 high-confidence results, including their associated contextual captions.
@@ 4.3. Compatibility
The CogPlanner framework is inherently agnostic to the specific model employed, making it easy to be integrated into any MRAG system and immediately enhancing their performance, demonstrating stunning flexibility in real-world applications. The planning expert responsible for query reformulation and determining the appropriate retrieval action can be any MLLM, or even a traditional classification model. For the planning expert, we exclusively employ a diverse set of advanced MLLMs as the foundation. Specifically, we leverage both closed-source APIs and open-source MLLMs. The closed-source models include GPT-4o, while the open-source models consist of the Qwen-VL series (Bai et al., 2023) and the Pixtral series (Dubey et al., 2024).
@@ 5. CogBench Construction
In this section, we present CogBench, a benchmark specifically developed for the MRAG planning task. CogBench comprises over 7,000 data samples, with a high-quality test set containing more than 400 samples. This benchmark is designed to facilitate the assessment of the effectiveness of our proposed CogPlanner framework, as well as other MRAG planning frameworks. Moreover, CogBench can be leveraged to enhance the decision-making capabilities of various MLLMs, particularly resource-efficient models, through fine-tuning. In the following subsections, we detail the construction process of CogBench and demonstrate how the benchmark enables lightweight integration of CogPlanner with the Qwen2-VL-7B-Instruct.
@@ 5.1. Query Collection
The rapid evolution of MLLMs has created an imperative need to evaluate these systems against increasingly sophisticated user queries that mirror real-world applications. While existing benchmarks (Jiang et al., 2024b; Hu et al., 2024) provide valuable groundwork, we recognized the necessity to extend beyond their scope. To this end, we deliberately incorporate more intricate queries that specifically demand image-based knowledge augmentation. We acquire authentic user interactions through web-sourced screenshots. We first curate a diverse array of topics and structure search queries around these topic words, such as ”CMU posters”. We use Google Image Search to collect image corpus. The resultant image corpus is subject to a stringent filtration process, based on a set of predefined quality criteria and a manual selection process. To generate realistic queries, we leveraged GPT-4o to simulate real users, producing five distinct questions per image that span both factual and open-ended inquiries requiring visual context interpretation. Each query-image pair undergoes manual review by two senior AI research engineers, each bringing at least three years of domain expertise. The engineers select the most compelling 1-5 queries per image that exemplify the strong potential for multimodal retrieval.
@@ 5.2. MRAG Planning and Generation
The MRAG planning and generation process is central to the CogPlanner framework, as detailed in Section 4.2. We employ GPT-4o API for executing the planning process. This implementation records each iteration of the planning process, encompassing the series of actions taken, the multimodal document sets retrieved, and the responses generated. Following response generation, the expert annotators conduct a thorough examination to review and regularize the entire chain of actions, and manually annotate the golden answer. Each data sample in the CogBench, therefore, contains a multimodal query, the retrieval actions, reformulated queries at each iterative step, the documents retrieved, and the final golden answer. The CogBench dataset is divided into training and test sets, comprising 5307 and 401 samples, respectively. A particularly noteworthy aspect of the annotation process is the possible variability in the decision chains for identical queries in the test set. While these chains may diverge considerably, the final reformulated queries, derived from the terminal pre-generation iteration, serve as a reflective ”result” of the entire planning procedure. These final queries encapsulate the cumulative informational requirements and explicit intent of the planning process, making it a crucial benchmark for assessing the accuracy of MRAG planning. Thus, the terminal queries are carefully constructed by human annotators.
@@ 5.3. CogBench Analysis
As shown in Table 1 and Figure 3, CogBench contains 5718 user queries spanning 9 distinct cognitive domains. We identify several fundamental characteristics that distinguish CogBench from existing benchmarks. (1) Unlike previous benchmarks, which only focus primarily on query-response pairs, CogBench offers a comprehensive record of the entire planning procedure involved in MRAG tasks, thereby facilitating the training of MLLMs. (2) A critical limitation of current MRAG benchmarks lies in their reliance on artificial query construction, primarily through simple entity substitution techniques (Lin and Byrne, 2022; Hu et al., 2024). Such methodologies typically yield responses confined to single entities or numerical values, severely understating the complexity inherent in real-world multimodal interaction scenarios. In contrast, CogBench introduces 24.19% open-ended queries that demand sophisticated, multi-faceted responses encompassing multiple interconnected claims with much longer answer length - 40.13 tokens on average. (3) CogBench incorporates diverse planning procedures that necessitate distinct search strategies at different stages, resulting in varied decision chains across different MLLMs. To quantify this complexity, we also ask the annotators to assess the number of reasoning steps required for resolution (e.g. a single-round of question answering is considered a 1-hop query). Our findings reveal that 79.55% cases explicitly require MRAG Planning, highlighting the sophisticated nature of the reasoning tasks presented in our benchmark.
@@ 5.4. Lightweight Integration of CogPlanner
To improve the efficiency and reduce the resource requirements of the CogPlanner framework, we aim to achieve a more lightweight integration within MRAG systems. Building upon the CogBench, we introduce a specialized fine-tuning strategy tailored for smaller, resource-efficient MLLMs to broaden the applicability of the CogPlanner framework and mitigate resource constraints. We utilize the Qwen2-VL-7B-Instruct as the backbone and employ the CogBench training set as specialized training data. To maintain a balance between the fine-tuning process and the retention of the Qwen2-VL-7B-Instruct model’s general capabilities, we augment the training dataset with general instruction data at a 1:1 ratio. This ensures that the model benefits from the specialized training required for CogPlanner integration while preserving its broad functionality. The fine-tuned model is referred to as Qwen2-7b-VL-Cog. This fine-tuning methodology is highly adaptable and capable of being applied to any existing MLLM. In our practice, we find that the CogBench fine-tuning process significantly enhances the MRAG planning capabilities of MLLMs, qualified for an effective planning expert. Ultimately, this approach facilitates the development of a lightweight integration of CogPlanner, enhancing MRAG performance while requiring minimal additional computational resources.
@@
Model Reasoning-Steps End-to-End Performance
@@
GPT-4o 0.3349 0.3744 0.1001 0.4438 0.5946 0.3984 0.1617 0.2474 0.1283 0.2907 0.3885 0.2121
Pixtral-Large-Instruct 0.2430 0.4187 0.0519 0.3765 0.5464 0.3451 0.1077 0.2450 0.0713 0.2245 0.3805 0.1581
QVQ-72B-Preview 0.3415 0.2210 0.1056 0.4051 0.3129 0.2148 0.1942 0.1399 0.0637 0.2943 0.2139 0.1224
Qwen2-VL-72B-Instruct 0.3445 0.4885 0.0795 0.3546 0.4494 0.3700 0.1255 0.1700 0.0927 0.2463 0.3278 0.1819
@@
GPT-4o 0.3173 0.2313 0.1074 0.3675 0.4514 0.3607 0.1788 0.301 0.1113 0.2697 0.3367 0.1932
Pixtral-Large-Instruct 0.2310 0.3927 0.0613 0.2805 0.4451 0.2928 0.1258 0.2491 0.0688 0.1986 0.3435 0.1416
QVQ-72B-Preview 0.1135 0.2313 0.0272 0.1710 0.1977 0.2042 0.0817 0.1024 0.0270 0.1178 0.1603 0.0858
Qwen2-VL-72B-Instruct 0.2630 0.3565 0.0902 0.2692 0.3244 0.2920 0.1645 0.1745 0.0789 0.2193 0.2614 0.1519
@@
GPT-4o 0.2268 0.1563 0.0572 0.3872 0.3679 0.3708 0.1572 0.1711 0.1049 0.2477 0.2333 0.1833
Pixtral-Large-Instruct 0.0835 0.1422 0.0157 0.1691 0.2173 0.2727 0.0858 0.1301 0.0718 0.1130 0.1615 0.1270
QVQ-72B-Preview 0.2922 0.1854 0.0804 0.2490 0.1662 0.1875 0.0995 0.0808 0.0602 0.1885 0.1305 0.1066
Qwen2-VL-72B-Instruct 0.2056 0.2447 0.0712 0.2666 0.2707 0.3064 0.1163 0.1333 0.0998 0.1844 0.2017 0.1625
@@
GPT-4o 0.3574 0.3767 0.1046 0.4746 0.5737 0.4283 0.2409 0.2773 0.1365 0.3422 0.3959 0.2267
Pixtral-Large-Instruct 0.2146 0.3708 0.0516 0.4535 0.5229 0.3896 0.2095 0.3034 0.1112 0.2915 0.3900 0.1913
QVQ-72B-Preview 0.3638 0.3644 0.4816 0.4390 0.2960 0.2443 0.2219 0.2522 0.2669 0.3229 0.2897 0.3033
Qwen2-VL-72B-Instruct 0.3550 0.4877 0.0823 0.4588 0.4708 0.3851 0.2164 0.2400 0.1571 0.3245 0.3690 0.2174
@@
GPT-4o 0.3292 0.3346 0.1014 0.4967 0.5468 0.4326 0.2803 0.3338 0.1526 0.3621 0.4046 0.2349
Pixtral-Large-Instruct 0.2260 0.3639 0.0557 0.3718 0.5432 0.3960 0.0996 0.2996 0.1128 0.2157 0.3936 0.1951
QVQ-72B-Preview 0.3562 0.4595 0.5169 0.4336 0.4407 0.2537 0.1954 0.2118 0.2727 0.3073 0.3384 0.3163
Qwen2-VL-72B-Instruct 0.3688 0.4886 0.0784 0.4294 0.4445 0.3748 0.2157 0.2501 0.1427 0.3179 0.3633 0.2065
@@ 6.1. Experimental Settings
Our experimental evaluation of CogPlanner, conducted on CogBench test set, encompasses two primary dimensions. First, we assess the comprehensive performance of the overall MRAG system to quantify CogPlanner’s effectiveness. Second, we conduct a focused analysis of CogPlanner’s planning procedure to validate our architectural design choices.
@@ 6.1.1. End-to-End MRAG Performance
We designed five distinct experimental configurations across all selected MLLMs. The baseline configuration employs the original MLLMs, where multimodal queries are processed directly by the MLLM. We then examine two intermediate configurations: one incorporating fixed image retrieval based on visual query components, and another utilizing fixed text retrieval driven by textual query components. The core evaluation focuses on both parallel and sequential modeling implementations of CogPlanner. For performance metrics, we adopt both token-level and claim-level evaluations, inspired by (Ru et al., 2024). Token-level evaluation is performed using the F1 score, measuring the overlap of common tokens between the generated response and the ground truth. Specifically, we use the NLTK tokenizer to segment the generated answers and ground truth. For claim-level evaluation, we utilize both precision and recall, calculated by first extracting claims from both the golden and generated answers using GPT-4o. Precision measures the proportion of correct claims within the generated responses, while recall evaluates the proportion of correct claims relative to the ground-truth answer claims.
6.1.2. Planning Procedure Performance
In addition to the overall MRAG performance, we examine the efficiency of CogPlanner’s planning procedure, with emphasis on the query reformulation action. We evaluate three distinct approaches: parallel modeling, sequential modeling, and direct reformulating query through GPT-4o with optimized prompt engineering. The comparative analysis employs standard metrics including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and F1 scores to assess the quality and relevance of the reformulated query outputs against established ground truth.
6.1.3. Implementation Details
We utilize the following backbone MLLMs, GPT-4o, QVQ-72B-Preview (Team, 2024), Qwen2-VL-72B-Instruct (Wang et al., 2024), Pixtral-Large-Instruct (Jiang et al., 2024a), and our fine-tuned Qwen2-7b-VL-Cog. In our retrieval process, we retain the top five results from a web search, each result containing a maximum of 800 tokens. For fine-tuning the Qwen2-7b-VL-Cog model, we leverage the Llama-Factory framework (Zheng et al., 2024) with a learning rate of 2e-6. A cosine learning rate scheduler is employed, and training proceeds for 2 epochs with a batch size of 32 and a warm-up ratio of 0.1. Our experiments are conducted on 8 NVIDIA A800 GPUs.
Table 3. Performance of the query reformulation task across different MLLMs.

### v2

@@ 4.2.2. Retrieval and Generation
At each iterative stage, the selection of retrieval actions determines whether or not and which retrieval API is invoked, either text retrieval or image retrieval. We leverage Google Web Search and Google Image Search as our primary retrieval API service. For each text retrieval request, we retrieve the top-k search results to ensure that only the most relevant information is retained. These results undergo preprocessing through the Jina API framework  [^2] https://github.com/jina-ai/reader, which transforms the raw web content into structured representations better suited for MLLM consumption. The visual retrieval pipeline captures full-page screenshots of search results, employing a set of systematic human-crafted rules to eliminate extraneous elements such as white space and original query images. To balance computational cost with retrieval quality, we limit the image retrieval to between three and six high-confidence candidates, accompanied by their contextual captions to provide relevant semantic grounding. Furthermore, to enhance system efficiency and responsiveness, we incorporate caching mechanisms for both text and image retrieval modules.
@@ 4.3. Compatibility
The CogPlanner framework is inherently agnostic to the specific model employed, making it easy to be integrated into any MRAG system and immediately enhancing their performance, demonstrating stunning flexibility in real-world applications. The planning expert responsible for query reformulation and determining the appropriate retrieval action can be any MLLM, or even a traditional classification model. For the planning expert, we exclusively employ a diverse set of advanced MLLMs as the foundation. Specifically, we leverage both closed-source APIs and open-source MLLMs. The closed-source models include GPT-4o (Hurst et al., 2024), while the open-source models consist of the Qwen-VL series (Bai et al., 2023) and the Pixtral series (Dubey et al., 2024).
@@ 5. CogBench Construction
In this section, we present CogBench, a benchmark specifically developed for the MRAG planning task. CogBench comprises over 5,000 data samples, with a high-quality test set containing more than 400 samples. This benchmark is designed to facilitate the assessment of the effectiveness of our proposed CogPlanner framework, as well as other MRAG planning frameworks. Moreover, CogBench can be leveraged to enhance the decision-making capabilities of various MLLMs, particularly resource-efficient models, through fine-tuning. In the following subsections, we detail the construction process of CogBench and demonstrate how the benchmark enables lightweight integration of CogPlanner with the Qwen2-VL-7B-Instruct.
@@ 5.1. Query Collection
The rapid evolution of MLLMs has underscored the need for evaluation on increasingly complex user queries that mirror real-world application scenarios. While existing benchmarks (Jiang et al., 2024b; Hu et al., 2024) provide valuable groundwork, we recognized the necessity to extend beyond their scope. To this end, we deliberately incorporate more complex queries that demand image-based knowledge augmentation. We acquire authentic user intent through web-sourced screenshots. We curate a diverse array of topics and structure search queries around these topic words, such as ”Astro Bot screenshot”, and use Google Image Search to collect an image corpus. The resultant image corpus is subject to a manual filtration process. To generate realistic queries, we leveraged the Claude-3.5-sonnet API  [^3] https://www.anthropic.com/news/claude-3-5-sonnet to simulate real users, producing five distinct queries per image that span both factual and open-ended inquiries requiring visual context interpretation. Each query-image pair undergoes manual review and modification by two senior AI research engineers, each bringing at least three years of domain expertise. The modification process follows several key principles: (1) each query must be distinct, with no repetition—even across different images; (2) queries must be unambiguous; (3) queries should be meaningful and formulated naturally, resembling how real humans would ask them; and (4) each query should target specific information, asking about concrete aspects of the image. For each image, the 1–5 most compelling queries, which highlight the potential of multimodal retrieval, are retained.
@@ 5.2. MRAG Planning and Generation
The MRAG planning and generation process is central to the CogPlanner framework, as detailed in Section 4.2. We employ the GPT-4o API for executing the planning process. This implementation records each iteration of the planning process, encompassing the series of actions taken, the multimodal document sets retrieved, and the responses generated. Following response generation, the expert annotators conduct a thorough examination to review and regularize the entire chain of actions, and manually annotate the golden answer. Each data sample in the CogBench, therefore, contains a multimodal query, the retrieval actions, reformulated queries at each iterative step, the documents retrieved, and the final golden answer. To be noticed, we do not define a fixed gold standard for the multimodal query processing, as manual annotation of information collection paths does not yield a unique or definitive reference. Instead, we focus on the correctness and completeness of the final answer. Finally, the CogBench dataset is divided into training and test sets, comprising 5307 and 401 samples, respectively.
@@ 5.3. CogBench Analysis
As shown in Table 1 and Figure 3, CogBench contains 5718 user queries spanning 9 distinct cognitive domains. We identify several fundamental characteristics that distinguish CogBench from existing benchmarks. (1) Unlike previous benchmarks, which focus primarily on query-response pairs, CogBench offers a comprehensive record of the entire planning procedure involved in MRAG tasks, thereby facilitating the training of MLLMs. (2) A critical limitation of current MRAG benchmarks lies in their reliance on artificial query construction, primarily through simple entity substitution techniques (Lin and Byrne, 2022; Hu et al., 2024). Such methodologies typically yield responses confined to single entities or numerical values, severely understating the complexity inherent in real-world multimodal interaction scenarios. In contrast, CogBench introduces 24.19% open-ended queries that demand sophisticated, multi-faceted responses encompassing multiple interconnected claims with much longer answer length - 40.13 tokens on average. (3) CogBench incorporates diverse planning procedures that necessitate distinct search strategies at different stages, resulting in varied decision chains across different MLLMs. To quantify this complexity, we also ask the annotators to assess the number of reasoning steps required for resolution (e.g. a single round of question answering is considered a 1-hop query). Our findings reveal that 79.55% of cases explicitly require MRAG Planning, highlighting the sophisticated nature of the reasoning tasks presented in our benchmark.
@@ 5.4. Lightweight Integration of CogPlanner
To improve the efficiency and reduce the resource requirements of the CogPlanner framework, we aim to achieve a more lightweight integration within MRAG systems. Building upon the CogBench, we introduce a specialized fine-tuning strategy tailored for smaller, resource-efficient MLLMs to broaden the applicability of the CogPlanner framework and mitigate resource constraints. We utilize the Qwen2-VL-7B-Instruct as the backbone and employ the CogBench training set as specialized training data. To maintain a balance between the fine-tuning process and the retention of the Qwen2-VL-7B-Instruct model’s general capabilities, we augment the training dataset with general instruction data at a 1:1 ratio. This ensures that the model benefits from the specialized training required for CogPlanner integration while preserving its broad functionality. The fine-tuned model is referred to as Qwen2-7B-VL-Cog. This fine-tuning methodology is highly adaptable and capable of being applied to any existing MLLM. In our practice, we find that the CogBench fine-tuning process significantly enhances the MRAG planning capabilities of MLLMs, making them qualified for effective planning experts. Ultimately, this approach facilitates the development of a lightweight integration of CogPlanner, enhancing MRAG performance while requiring minimal additional computational resources.
@@
Model Reasoning-Steps Overall Performance
@@
GPT-4o 33.49 37.44 10.01 44.38 59.46 39.84 16.17 24.74 12.83 29.07 38.85 21.21
Pixtral-Large-Instruct 24.30 41.87 5.19 37.65 54.64 34.51 10.77 24.50 7.13 22.45 38.05 15.81
QVQ-72B-Preview 34.15 22.10 10.56 40.51 31.29 21.48 19.42 13.99 6.37 29.43 21.39 12.24
Qwen2-VL-72B-Instruct 34.45 48.85 7.95 35.46 44.94 37.00 12.55 17.00 9.27 24.63 32.78 18.19
@@
GPT-4o 31.73 23.13 10.74 36.75 45.14 36.07 17.88 30.10 11.13 26.97 33.67 19.32
Pixtral-Large-Instruct 23.10 39.27 6.13 28.05 44.51 29.28 12.58 24.91 6.88 19.86 34.35 14.16
QVQ-72B-Preview 11.35 23.13 2.72 17.10 19.77 20.42 8.17 10.24 2.70 11.78 16.03 8.58
Qwen2-VL-72B-Instruct 26.30 35.65 9.02 26.92 32.44 29.20 16.45 17.45 7.89 21.93 26.14 15.19
@@
GPT-4o 22.68 15.63 5.72 38.72 36.79 37.08 15.72 17.11 10.49 24.77 23.33 18.33
Pixtral-Large-Instruct 8.35 14.22 1.57 16.91 21.73 27.27 8.58 13.01 7.18 11.30 16.15 12.70
QVQ-72B-Preview 29.22 18.54 8.04 24.90 16.62 18.75 9.95 8.08 6.02 18.85 13.05 10.66
Qwen2-VL-72B-Instruct 20.56 24.47 7.12 26.66 27.07 30.64 11.63 13.33 9.98 18.44 20.17 16.25
Self-Reflective RAG
GPT-4o 43.83 11.72 15.56 28.79 22.21 30.88 22.79 20.68 13.03 29.59 19.13 19.47
Qwen2-VL-72B-Instruct 43.29 10.97 10.49 27.55 20.72 28.58 19.01 18.96 13.85 26.81 17.91 18.05
@@
GPT-4o 35.74 37.67 10.46 47.46 57.37 42.83 24.09 27.73 13.65 34.22 39.59 22.67
Pixtral-Large-Instruct 21.46 37.08 5.16 45.35 52.29 38.96 20.95 30.34 11.12 29.15 39.00 19.13
QVQ-72B-Preview 36.38 36.44 48.16 43.90 29.60 24.43 22.19 25.22 26.69 32.29 28.97 30.33
Qwen2-VL-72B-Instruct 35.50 48.77 8.23 45.88 47.08 38.51 21.64 24.00 15.71 32.45 36.90 21.74
@@
GPT-4o 32.92 33.46 10.14 49.67 54.68 43.26 28.03 33.38 15.26 36.21 40.46 23.49
Pixtral-Large-Instruct 22.60 36.39 5.57 37.18 54.32 39.60 9.96 29.96 11.28 21.57 39.36 19.51
QVQ-72B-Preview 35.62 45.95 51.69 43.36 44.07 25.37 19.54 21.18 27.27 30.73 33.84 31.63
Qwen2-VL-72B-Instruct 36.88 48.86 7.84 42.94 44.45 37.48 21.57 25.01 14.27 31.79 36.33 20.65
@@ 6.1. Experimental Settings
Our experimental evaluation of CogPlanner, conducted on the CogBench test set, encompasses two primary dimensions: the overall performance of the MARG system and the analysis of the planning procedure within CogPlanner. We utilize the following backbone MLLMs, GPT-4o (Hurst et al., 2024), Qwen2-VL-72B-Instruct (Wang et al., 2024), Pixtral-Large-Instruct (Jiang et al., 2024a), and our fine-tuned Qwen2-7B-VL-Cog. Notably, QVQ-72B-Preview serves as a representative MLLM for advanced multimodal reasoning capabilities.
@@ 6.1.1. End-to-End MRAG Performance
We conduct six distinct experimental configurations across all selected MLLMs. The baseline configuration employs the original MLLMs, where multimodal queries are processed directly by the MLLM. We then examine two intermediate configurations: one incorporating fixed image retrieval based on visual query components, and another utilizing fixed text retrieval driven by textual query components. Besides, we employ the self-reflective RAG framework [^4] https://github.com/langchain-ai/langgraph/tree/main/examples/rag (Asai et al., 2023) as a reflective and iterative competitive framework. The core evaluation focuses on both parallel and sequential modeling implementations of CogPlanner.
For performance metrics, we adopt both token-level and claim-level evaluations, inspired by (Ru et al., 2024). Token-level evaluation is performed using the F1 score, measuring the overlap of common tokens between the generated response and the ground truth. Specifically, we use the NLTK tokenizer to segment the generated answers and ground truth. For claim-level evaluation, we utilize both precision and recall, calculated by first extracting claims from both the golden and generated answers using GPT-4o. Precision measures the proportion of correct claims within the generated responses, while recall evaluates the proportion of correct claims relative to the ground-truth answer claims.
Table 3. Performance of query reformulation across different MLLMs.
@@
6.1.2. Planning Procedure Performance
In addition to the overall MRAG performance, we examine the efficiency of CogPlanner’s planning procedure, with emphasis on query reformulation, by comparing its reformulated queries with those annotated by human experts. We evaluate three distinct approaches: parallel modeling, sequential modeling, and direct reformulating query through GPT-4o with optimized prompt engineering. The comparative analysis employs standard metrics, including BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and F1 scores to assess the quality and relevance of the reformulated query outputs against established ground truth.
Table 4. End-to-end performance and efficiency evaluation of Qwen2-VL-72B-Instruct and Qwen2-7B-VL-Cog as planning experts, with response generation models consistently using Qwen2-VL-72B-Instruct.
Planning Expert Precision Recall F1 Avg # Total Tokens Latency(s)
CogPlanner With Parallel Modeling
Qwen2-VL-72B-Instruct 32.45 36.90 21.74 30.36 9.76 (14.9%) 1.209
Qwen2-7B-VL-Cog 31.97 32.65 21.46 28.69 7.58 (9.8%) 0.484
CogPlanner With Sequential Modeling
Qwen2-VL-72B-Instruct 31.79 36.33 20.65 29.59 13.56 (21.6%) 1.842
Qwen2-7B-VL-Cog 32.50 33.10 21.38 29.00 7.59 (11.9%) 0.545
6.1.3. Implementation Details
In our retrieval process, we retain the top five results from a web search, each result containing a maximum of 800 tokens. For fine-tuning the Qwen2-7B-VL-Cog model, we leverage the Llama-Factory framework (Zheng et al., 2024) with a learning rate of 2e-6. A cosine learning rate scheduler is employed, and training proceeds for 2 epochs with a batch size of 32 and a warm-up ratio of 0.1. To ensure computational efficiency and prevent indefinite reasoning loops, we impose an upper bound of three iterations on the CogPlanner planning process. Our experiments are conducted on 8 NVIDIA A800 GPUs.